### v1

# Indoor Location Fingerprinting Privacy: A Comprehensive Survey

Abstract.
The pervasive integration of Indoor Positioning Systems (IPS) arises from the limitations of Global Navigation Satellite Systems (GNSS) in indoor environments, leading to the widespread adoption of Location-Based Services (LBS). Specifically, indoor location fingerprinting employs diverse signal fingerprints from user devices, enabling precise location identification by Location Service Providers (LSP). Despite its broad applications across various domains, indoor location fingerprinting introduces a notable privacy risk, as both LSP and potential adversaries inherently have access to this sensitive information, compromising users’ privacy. Consequently, concerns regarding privacy vulnerabilities in this context necessitate a focused exploration of privacy-preserving mechanisms. In response to these concerns, this survey presents a comprehensive review of Privacy-Preserving Mechanisms in Indoor Location Fingerprinting (ILFPPM) based on cryptographic, anonymization, differential privacy (DP), and federated learning (FL) techniques. We also propose a distinctive and novel grouping of privacy vulnerabilities, adversary and attack models, and available evaluation metrics specific to indoor location fingerprinting systems. Given the identified limitations and research gaps in this survey, we highlight numerous prospective opportunities for future investigation, aiming to motivate researchers interested in advancing this field. This survey serves as a valuable reference for researchers and provides a clear overview for those beyond this specific research domain.
Indoor Localization, Privacy-Preserving, Location Fingerprinting
ccs: Security and privacy Privacy protections
1. Introduction
Localization is the process of determining the spatial coordinates or position of a tracked item with respect to several reference points within a predefined area (Geng et al., 2021). There are growing appeals for providing Location-based services (LBS) to people around the world. The appearance of Global Navigation Satellite Systems (GNSS) was a turning point in localization, but lately, the traditional scope of positioning systems evolved with the advent of mobile computing, including the Internet of Things (IoT) and wearable devices, moving from military tracking and civilian navigation to location information. For both personal and commercial purposes, location data is now a crucial component in bridging the gap between the real and digital worlds. Nowadays, due to the widespread use of mobile devices, users are constantly generating various location data during their daily activities. Accordingly, there is a growing interest in LBS, which offer services depending on the user’s location. Such services assist people in finding the appropriate route to different destinations, selecting various points of interest (POI) based on their preferences, and even remembering to take precautions for their healthcare purposes (Pascacio et al., 2021).
All these LBSs require reliable and real-time location information, which is often accomplished using GNSS technologies. However, due to the severe attenuation of satellite signals and the low visibility of satellites in indoor environments, they mostly fail for indoor localization purposes (Asaad and Maghdid, 2022). The main indoor localization technologies are alternative signals such as Wi-Fi, Bluetooth, FM, AM, GSM or LTE (Moghtadaiee and Dempster, 2014; Rahman et al., 2017). During localization, they all should face indoor area challenges such as non-line-of-sight (NLOS) problems, multipath propagation, signal blockage, intentional and unintentional interferences, etc.
Indoor positioning has gotten so much attention in the last decade as individuals spend 80% of their time indoors (Kepeis et al., 2001), which shows the significant role of indoor localization. Unlike in outdoor positioning systems, the user has no direct access to her location, and it is calculated by the Location Service Provider (LSP), the server in the Indoor Positioning System (IPS), based on the signal information received from the user, as shown in Fig. 1. Moreover, more accuracy is required for indoor localization than it is for outdoor localization which only needs meter-level accuracy. The indoor environments are diverse and include residences, workplaces, shopping malls, warehouses, hospitals, and retail centers. Additionally, there are many various applications for end users with different requirements for accuracy and coverage. For instance, while law-enforcement applications with urban/rural coverage have an accuracy of a few meters, ambient assisted-living applications require room-level coverage with an accuracy below $1m$ (Mautz, 2012). Therefore, no IPS has evolved as an all-encompassing solution due to the unique characteristics of indoor settings and the diversity of applications. Not only the coverage and accuracy are important to consider when selecting the base positioning technology, but also the deployment and maintenance costs are crucial. The diverse solutions are a clear indicator that there is no single alternative for GNSS in indoor areas, and different positioning technologies co-exist.
Figure 1. Outdoor vs. indoor positioning systems.
However, with all the benefits provided by IPS, users are deeply concerned about their privacy when their personal location data is accessed; hence, proper privacy preservation is of utmost importance for indoor LBS applications, motivated many researchers to work on how to protect the privacy of user locations in indoor environments. The development of localization methodologies has also made it possible for indoor LBS providers to quickly gather a significant amount of location data from several users, which is necessary to better understand user behavior and, as a result, improve the quality of the services, but it severely violates the privacy of users at the same time.
In response to privacy concerns, many countries have enacted privacy laws and regulations, such as the General Data Protection Regulation (GDPR) [^1] https://gdpr-info.eu/ in the European Union, and the California Consumer Privacy Act (CCPA) [^2] https://oag.ca.gov/privacy/ccpa in the United States, to provide individuals with greater control over their personal data, including location information. Specifically for location privacy, an architecture of an indoor localization model ensuring GDPR compliance through the integration of a specialized GDPR-based access control system is introduced in (Barsocchi et al., 2020).
Therefore, the main obstacle to the widespread adoption of IPS in the world is privacy. When a user is linked to the network, IPS always ”knows” where they are. Without the user’s prior consent, tracking her location is prohibited as users are generally reluctant to share their whereabouts. Location data reveal private and sensitive information on each user, including their health conditions, their interests, their views and behaviors, their usage amount of electricity, their workplaces, and their home and job locations. The leakage of private assembly line information, sensitive process information, and other relevant information are additional security risks posed by IPS in commercial or high-profile buildings. Moreover, the radio map of the buildings can be used to access the structural configuration, which might raise terrorist attacks and other negative actions. Therefore, practical answers are required to these privacy problems (Singh et al., 2021). Authorities will establish stringent but pertinent norms and regulations for industries to use users’ location data as nontechnical solutions. Without the user’s explicit consent, data should not be shared with third-party organizations or people. Therefore, the lack of comprehensive efforts on this topic motivates us to examine indoor location privacy from various points of view thoroughly.
Several papers are surveying IPS in terms of accuracy, localization techniques, and technologies (Singh et al., 2021; Nessa et al., 2020; Zafari et al., 2019; Ngamakeur et al., 2020). Considering the security and privacy of IPS, on the other hand, there are two limited surveys (Holcer et al., 2020; YILDIZ and PEKEY, 2019) and only one comprehensive survey (Sartayeva and Chan, 2023).
The first review paper in 2020, (Holcer et al., 2020), conducts a limited analysis of privacy in IPS using the Preferred Reporting Items for Systematic Review and Meta-Analysis (PRISMA) guidelines (Liberati et al., 2009). The paper categorizes the existing research papers (up to 2020) into three groups: privacy on the device, during transmission, and at the server, and explains them based on this categorization. However, as this research is from 2020, the research studies after that time are not included. In 2022, the only comprehensive survey (Sartayeva and Chan, 2023) categorizes the indoor positioning methods into two items: (1) non-collaborative methods including proximity-based methods, geometry methods, fingerprinting, and others, and (2) collaborative methods including mobile proximity-based methods, mobile geometric methods, and others. For the sake of privacy and security, the survey applies a confidentiality, integrity, and authenticity trilogy to each of the collaborative and non-collaborative methods. The main focus of this paper, however, is on the security issues of IPS. Regarding various privacy attacks and privacy-preserving techniques when location fingerprinting technique is used, it only mentions limited number of publications. The last existing survey in 2023, (YILDIZ and PEKEY, 2019), first investigates the existing reviews on IPS and the technologies and techniques applied to them, followed by their advantages and disadvantages. It explains some of the major challenges for IPS. In terms of privacy in indoor location fingerprinting, it surveys a restricted number of studies, 15 papers between 2018 and 2022, without employing categorization, most of which are based on secure two-party computation and homomorphic encryption. Furthermore, it neither takes into account more recent papers introducing new methodologies nor discusses privacy attacks.
This paper focuses on privacy-preserving mechanisms in the context of indoor location fingerprinting, the most widely used localization technique in IPS. Throughout the rest of the paper, when referring to IPS, we specifically mean IPS utilizing location fingerprinting as its localization technique. The contributions of this survey paper are outlined as follows:
  • To the best of our knowledge, this is the first deep dive survey considering indoor location fingerprinting privacy-preserving mechanisms (ILFPPM) from various novel aspects: indoor location privacy definitions, applications, entities vulnerabilities, possible attacks, privacy protection techniques, challenges, privacy metrics, and vision of future research directions.
  • In addition to discussing IPS fundamentals, our exploration includes an thorough analysis of the various applications of IPS, investigating their practical implementations and addressing the privacy concerns associated with the collection, storage, and utilization of location data in indoor spaces.
  • We comprehensively identify privacy vulnerabilities and all sources of privacy leakages in indoor location fingerprinting systems, examining them from diverse perspectives encompassing system entities, indoor location data structure, and information inference. Additionally, we introduce a novel categorization of adversary models and attack models specific to these systems, marking the first systematic classification in this domain
  • To facilitate empirical investigations into ILFPPM, we provide list of datasets from prior research. We also categorize, for the first time, all metrics employed in existing studies. This endeavor is intended to establish a robust benchmark within the research community, fostering future empirical analyses and enhancing technical insights into this domain.
  • Through the literature analysis, we have outlined the unresolved challenges and suggested various potential paths for future research in location privacy attacks and ILFPPM. Our aim is to motivate readers with an interest in this field to further explore these issues.
As outlined in Fig. 2, the rest of the paper is organized as follows. Section 2 presents IPS fundamentals including methodology, techniques, and technologies. Section 3 explains applications of IPS with their privacy concerns. Section 4 discusses the sources of privacy leakages in indoor location fingerprinting, followed by the adversary and attack models in Section 5. The existing ILFPPM are fully reported in Section 6. All the datasets and metrics utilized in these studies also presented in this Section. Section 7 discusses the future research directions, and Section 8 concludes the paper.
Figure 2. Taxonomy of this paper.
2. Indoor Positioning System Fundamentals
To understand privacy protection methods, it is important to first understand the fundamentals of IPS. IPS is a GPS-free system that continuously predicts the positions of people or objects in indoor environments by first applying a distance estimation algorithm, followed by the localization algorithm (Liu et al., 2007). Using a suitable ranging technique, an IPS calculates the distance between the target and anchor nodes whose coordinates are known a priori during the distance measurement phase. Using these distance observations, the IPS then uses several localization/positioning techniques to approximate the target’s location.
2.1. Ranging Techniques
Here we discuss various ranging techniques that have been employed for indoor localization so far.
2.1.1. Received Signal Strength (RSS)
In general, Received Signal Strength (RSS) represents the intensity of a signal measured at the receiver, typically at a certain distance from a signal source. It serves as one of the most straightforward signal metrics to measure, providing valuable insights into the strength of wireless signals. However, particularly in indoor environments, RSS measurements can be susceptible to various factors that introduce inaccuracies. These factors include fading, shadowing, refraction, scattering, and reflections within complex indoor layouts.
Due to the sensitivity of RSS measurements to these environmental influences, they often yield erroneous distance estimates in IPS. To mitigate the fluctuations and inaccuracies associated with RSS, various techniques and machine learning (ML) approaches have been adopted (Ben Kilani et al., 2014). These methods aim to enhance the reliability and precision of RSS-based positioning. Some of the strategies employed include the implementation of sophisticated filters and the utilization of ML algorithms. These approaches are designed to filter out noise, account for signal variations, and extract meaningful information from RSS data, ultimately improving the performance and robustness of IPS.
When measuring the RSS fingerprinting values is not feasible in a specific environment, or a very emergency training procedure is required, one way is to generate fingerprinting RSS values instead. For generating RSS values in an indoor area, the long-distance path-loss model has been frequently utilized, as it demonstrates how RSS measurements vary with distance from an AP (Chen and Terzis, 2011). The RSS values sensed from the $j$’th AP, ${S}_{u,j}$(dBm) at a physical distance $d_{j}$ in meter, are calculated as follows:
(1) ${RSS}_{u,j}={s_{d_{0}}}_{j}-10\alpha_{j}\log\left(\frac{d_{j}}{d_{0}}\right)+X_{\sigma_{j}},$
where ${s_{d_{0}}}_{j}$dBm is the power received at $d_{0}$m (usually $1$m) from $j$th AP, and $\alpha_{j}$ is the path-loss exponent associated with that AP. Due to the shadowing in indoor environments, $X_{\sigma_{j}}\sim\mathcal{N}(0,\sigma_{j})$ models path-loss variation at a single point and is assumed to be a zero mean Gaussian random variable with a standard deviation given by $\sigma_{j}$dB.
2.1.2. Signal Propagation Time
The distance between the target and the anchor node is also calculated based on the signal propagation time. These methods are typically more accurate than the RSS method. Time of Arrival (ToA) and Time Difference of Arrival (TDoA) are the two well-known methods of distance calculation. ToA-based algorithms require synchronizing between nodes, and TDoA methods were proposed to tackle it. Due to the fact that TDoA takes into account the synchronization of the transmitters, it can somewhat address the problem of synchronization inaccuracy (Do and Yoo, 2014). However, the performance of the ToA/TDoA-based systems is considerably hampered by the NLOS propagation of the signal (Guvenc et al., 2007).
2.1.3. Phase and Angle of Arrival
The Angle of Arrival (AoA) and Phase of Arrival (PoA) approach estimate position by utilizing the angle and the phase that a signal makes with an antenna array (Wielandt et al., 2018). This improved ranging method is used to measure both the angle and the distance. The need for antenna arrays, which makes this approach complex and expensive, is one of its limitations (Sakpere et al., 2017). The signal’s timing of arrival at certain antenna elements may also be used in this approach, although this requires much more complicated hardware and precise calibration.
2.1.4. Channel State Information
Channel State Information (CSI) is another advanced method for calculating distances in indoor positioning. Unlike RSS, which provides only the amplitude value of a received signal, CSI offers a more precise estimation of the received signal’s characteristics across the entire signal bandwidth. This improvement in accuracy is achieved by capturing the channel’s frequency response received by each antenna, making CSI a valuable tool for distance estimation. While CSI often requires multiple antennas to gather this information effectively, it can be employed in a variety of localization systems, including both range-based and range-free approaches (Yang et al., 2013). One notable advantage of CSI is its ability to provide not only signal amplitude but also phase information, enhancing the richness of data available for positioning calculations.”
2.2. Localization Methods
The localization methods which are commonly used for indoor localization are listed below:
2.2.1. Multilateration and Trilateration
With the aid of three or more known nodes and the corresponding associated distances, it is a method for determining the position of the unknown node (Carotenuto et al., 2019). Only three well-known nodes are employed in trilateration, which is a specific instance of multilateration. The target node’s location in a two-dimensional space is determined by the intersection of three fictitious circles. However, because of the NLOS effect, which results in significant positioning errors, these circles do not converge at a single location in the real-world indoor environment.
2.2.2. Triangulation
When the angle of arrival is known, it can be applied for positioning precision. It requires at least two anchor nodes and is moderately precise while being less complex (Pomárico-Franquiz et al., 2014). The accuracy of the AoA estimation is crucial to the technique’s location accuracy. The performance of localization can be improved by increasing the number of anchor nodes.
2.2.3. Fingerprinting
Fingerprinting is a widely used indoor location technique that utilizes Wi-Fi, BLE, FM, AM, and ZigBee wireless access technologies (Yan et al., 2018; Uradzinski et al., 2017; Moghtadaiee and Dempster, 2014; Rahman et al., 2017; Faragher and Harle, 2015). As shown in Fig. 3, two phases are involved in the fingerprint-based localization method: offline and online. A radio map is created using the measured data for each recorded position during the training phase using RSS or CSI data acquired at access points (APs) for various known indoor points called reference points (RPs). The Received Signal Strength (RSS) values from all existing APs at known Reference Points (RPs), along with the corresponding $(x,y)$ coordinates of RPs, are stored in the radiomap within LSP. The recorded radiomap is characterized by a dimension of $M\times N$, where $M$ and $N$ represent the number of RPs and APs, respectively. The fingerprint at a given location $(x_{i},y_{i})$ is denoted as ${rss}_{i}^{RP}=[{rss}_{i1},{rss}_{i2},\ldots,{rss}_{iN}]$, where ${rss}_{ij}$ represents the RSS of ${AP}_{j}$ at the ith RP. Moving to the online phase, the user’s location fingerprint is obtained as a vector of RSS values, $S_{u}=[{RSS}_{u1},{RSS}_{u2},\ldots,{RSS}_{uN}]$, where ${RSS}_{uj}$ signifies the RSS of $AP_{j}$ received by the user. This vector is then transmitted to LSP for localization purposes. During this phase, the user’s RSS vector is compared with the previously stored RSS values in the radiomap, and using machine learning methods, LSP estimates the user’s position, $(x,y)$. If more accurately obtained offline data is used to build the radio map, this method offers great accuracy. However, it takes a great deal of work to create the radio map for large-area deployment (e.g., manpower, time, and cost). Additionally, for dynamic networks, the offline database needs to be regenerated whenever a node’s position, even a single node, is unexpectedly changed or removed.
Figure 3. The flow of indoor location fingerprinting technique.
ML algorithms are frequently used to improve fingerprinting accuracy and aid in building radio maps (Nessa et al., 2020). The simplest algorithm used for fingerprint-based localization techniques is k-Nearest Neighbors (kNN). This algorithm computes a distance metric that measures the separations between the target’s measurements at various APs and the measurements taken during the training phase. Euclidean distance is the most widely used measurement of distance. The radio map’s k closest RPs to the target with the smallest distances are chosen in this algorithm. The target’s position is then inferred by averaging these RPs’ coordinates. However, the kNN algorithm also uses alternative distance measures such as the Manhattan, Mahalanobis, and Minkowski distances (Farshad et al., 2013).
2.2.4. Range-free
Instead of calculating the position based on distance, range-free localization methods are proposed:
  • Centroid: Instead of using distance or angle measurement, this method estimates the location of the unknown node using a geometric relation. Once a reliable communication link has been established between each anchor node and the unknown target node, the positions of the anchor nodes are known. The centroid of the geometric shape formed by the position of the anchor nodes connected to the target node is taken to be the location of the unknown nodes. In (Chen et al., 2013), a Weighted Centroid Localization (WCL) approach-based BLE beacon IPS was suggested.
  • Distance Vector Hop (DV Hop): In a multi-hop setting, this method involves estimating the distance vector depending on the hop count. An information table is kept up to date with the coordinates of the ith node and the minimum hop count value from the anchor node to the ith node. The anchor node broadcasts the position data to the nearby nodes, which subsequently broadcast it to more nodes, and so on. Finding the hop size for a specific hop is a crucial issue for this strategy. The distance between the node that is m hops distant from the anchor node is easily estimated after obtaining the average hop size (Chen et al., 2013).
2.3. Technologies
In this subsection, radio frequency wireless technologies that are most commonly used in indoor localization are briefly presented. A diverse array of wireless technologies is employed to enable precise tracking and positioning of objects, devices, or individuals within indoor environments. The choice of technology depends on factors such as accuracy requirements, infrastructure availability, power efficiency, and the specific application domain. In this subsection, two prominent categories of such technologies which are radio frequency (RF) and non-radio frequency wireless technologies are briefly discussed.
2.3.1. Radio Frequency
Radio frequency (RF) technologies are frequently deployed for indoor localization due to their versatility and robust performance in various settings. Among the widely utilized RF technologies are Wi-Fi-based localization systems, which capitalize on the ubiquity of Wi-Fi access points in indoor spaces. These systems employ signal strength and triangulation techniques to estimate the position of Wi-Fi-enabled devices such as smartphones or tablets (Yang and Shao, 2015). Bluetooth Low Energy (BLE) beacons also play a crucial role, especially in asset tracking and proximity-based applications. BLE beacons emit low-power signals that can be detected by smartphones or dedicated receivers, facilitating accurate indoor positioning (Faragher and Harle, 2014). Ultra-wideband (UWB) technology, recognized for its high precision, is gaining traction in applications demanding centimeter-level accuracy, such as industrial robotics and healthcare (Alarifi et al., 2016). FM/AM infrastructures, from another perspective, can be used to deploy various indoor localization methods without the need for additional hardware, which is applicable for long-range solutions with susceptibility to interference  (Moghtadaiee and Dempster, 2014; Rahman et al., 2017).
2.3.2. Non-Radio Frequency
Conversely, non-radio frequency wireless technologies are emerging as alternatives to RF solutions. Infrared (IR) technology relies on the transmission of light signals in the infrared spectrum to determine the location of objects equipped with IR transmitters and receivers. Although less common than RF technologies, IR systems offer advantages in specific scenarios, such as indoor positioning within controlled environments or line-of-sight requirements. Additionally, visible light communication (VLC) leverages LED lighting infrastructure to transmit data and positional information through modulated light signals. VLC systems have demonstrated potential for indoor localization while providing energy-efficient illumination (Alam et al., 2020).
3. Indoor Localization Applications and Privacy Concerns
In recent years, indoor localization advances and the proliferation of mobile devices have made LBS more accessible. The applications of indoor localization include but are not limited to:
3.1. Marketing
Contextually-aware location-based marketing strategies have the potential to significantly boost sales and profits in e-commerce by delivering tailored marketing directly to consumers based on their location, especially in shopping centers, utilizing location-based technology. This approach allows real-time communication between consumers and sellers, enhancing the overall shopping experience. Additionally, it aids businesses in tracking customer behaviors, patterns, and foot traffic, offering personalized services based on mobile devices’ information (Nessa et al., 2020; Singh et al., 2021). However, these strategies bring privacy challenges, as the use of location-based technology for personalized marketing raises concerns about user privacy and potential misuse of real-time location data. Tracking customer behaviors without explicit consent and the extensive collection of personal data through personalized mobile devices also pose ethical and privacy issues. Furthermore, the precision of IPS in localizing mobile devices indoors and outdoors intensifies concerns about intrusive tracking if not appropriately safeguarded.
3.2. Ambient Assisted Living and Disaster Management
Ambient assisted living platforms, relying on precise indoor location tracking through technologies like Bluetooth and various IPS, offer significant benefits for the elderly, ill, or disabled individuals, particularly those with neurodegenerative diseases (Wan et al., 2015). These platforms facilitate behavioral tracking, monitoring daily activities, movement patterns, vital signs, and detecting potential dangers such as falls or injuries. Smart houses leverage indoor localization to enhance user experiences, enabling the homeowner to control Wi-Fi network access based on device presence (Bonafini et al., 2019). Beyond homes, location tracking techniques find applications in agriculture for monitoring greenhouses, libraries for book location, parking garages for car monitoring, and warehouses for item tracking (Singh et al., 2021). In disaster management, indoor localization can pinpoint the precise location of individuals in peril for efficient rescue efforts (Chehri et al., 2011; Asimakopoulou and Bessis, 2011). Despite these advantages, privacy concerns arise, including the continuous monitoring of sensitive health-related data, potential invasions of privacy in smart homes, perceived intrusive surveillance in various contexts, and privacy implications in disaster management. Striking a balance between the benefits and privacy protection is crucial, necessitating robust data safeguards and ethical guidelines.
3.3. Health Services and Public Safety
Indoor localization holds immense potential for elevating service standards in healthcare, aiding hospital staff in swiftly locating patients and enabling patients to navigate therapy rooms independently. Doctors can monitor patient mobility, visitors can easily find their patients, and operating rooms can be well-equipped (Singh et al., 2021; Calderoni et al., 2015). Smart devices’ positions are traceable, facilitating their quick retrieval. The technology’s future applications include nanosensor-based drug delivery for tumor targeting and hazard identification for disaster mitigation (Giuliano et al., 2013). In law enforcement, Bluetooth beacon-based systems aid quick responses, while in smart buildings, Wi-Fi-enabled alarms guide people to safety. Despite these benefits, concerns arise about continuous tracking compromising patient privacy in healthcare, nanosensor-based drug delivery raising intimate monitoring questions, and surveillance implications in safety management and law enforcement. Striking a balance requires careful ethical and legal considerations in these applications.
During the COVID-19 pandemic, for example, IPS were employed to monitor and enforce social distancing guidelines in various settings, including hospitals and workplaces. While crucial for public health, these technologies raise significant privacy concerns. The continuous and precise tracking of individuals’ movements within indoor spaces may result in the collection of sensitive location data, revealing interactions, routines, and patterns. This raises potential privacy infringements, and there’s a risk of unauthorized access or misuse, challenging data security and individual autonomy. Addressing this complex challenge requires robust privacy safeguards, transparent data handling practices, and clear regulations to ensure the responsible deployment of IPS for COVID-19 measures while respecting privacy rights and maintaining public trust.
3.4. Surveillance and Tracking
LBS extend beyond human tracking, finding significant applications in industrial settings, where the automatic tracking of numerous objects is essential for effective management. This involves real-time control of item locations alongside localization and identification, requiring new Medium Access Control layer protocols to prevent collisions (Li, 2021). The prediction of autonomous robot locations, especially in the emerging era of collaborative robots (co-bots), enhances safety and efficiency in various tasks, such as safety and collision avoidance in industrial duties (Bentley et al., 2014; Chekuri and Won, 2017; Sadowski and Spachos, 2018). However, the continuous monitoring of objects and devices in industrial contexts raises privacy concerns for workers, impacting workplace privacy. Anticipating robot locations and monitoring employee interactions could similarly affect privacy. Furthermore, the use of location-based access control and continuous tracking for anti-theft systems raises data security and surveillance issues. Additionally, indoor localization in entertainment and broadcasting may involve tracking event attendees, presenting privacy challenges. Achieving a balance between the advantages of these technologies and privacy protection requires the establishment of robust safeguards and guidelines for responsible use.
4. Privacy Leakage In Indoor Location Fingerprinting
This section provides an in-depth examination of the indoor location fingerprinting system emphasizing privacy vulnerabilities and leakages. We provide a comprehensive understanding of the system entities and the underlying data structure, particularly in the context of employing location fingerprinting techniques for indoor localization. Through this, we derive insights into potential areas of privacy leakage within these systems. Additionally, we extend this section to a detailed exploration of the broader privacy leakage arising from the information inferred about the users in indoor location fingerprinting.
4.1. Entities-Driven Privacy Leakage
Here, we introduce and analyze the fundamental components associated with the processes of data transmission and service provisioning in the indoor location fingerprinting system so that we can comprehend the potential privacy vulnerabilities. As shown in Fig. 4, the entities include Access Points (APs), Users, Trusted Third Parties (TTPs), LBS servers, Content Providers, and Data Publishers. For each entity, we first introduce their characteristics, followed by the inherent privacy leakages within each entity, and then the potential privacy vulnerabilities that these entities may introduce.
Figure 4. Privacy-vulnerable entities in indoor location fingerprinting systems.
4.1.1. Access Point (AP)
The first hop in a data transmission often involves communication networks, mostly wireless APs. Wi-Fi, BLE, UWB, and RFID are among the applicable APs for indoor localization and data transmission. APs facilitate the transmission and reception of wireless signals within their coverage area. Devices seeking location information, such as smartphones or IoT devices, communicate with these APs through wireless signals. APs provide the means for wireless devices to connect to a network, enabling communication between the devices and the central network or server. This connectivity is essential for the exchange of location-related data during the localization process.
The potential privacy leakage for APs are the disclosure of their locations. The transmission of RSS/CSI data by APs to users during the service stage can potentially expose APs’ locations through various methods. The disclosure of this data for APs, particularly through RSS, makes them susceptible to malicious activities, leading to physical access and potential damage. The distance from an AP can be calculated using RSS information, and multilateration and triangulation techniques can be applied to pinpoint the AP’s location. Researchers have also found that the ratio of signal powers between the first peak and second peak in CSI can be used as a physical layer metric to estimate the distance from the AP (Wang et al., 2015).
APs can be also curious and pose a privacy threat themselves. They are indispensable in IPS, but they have less power than a malicious localization server. The user collects one measuring value from each AP during localization. Theoretically, localization requires at least three measuring values for a 2-dimensional area, (or four for a 3D area) which are located in different places. If the communication during measurement is unilateral, the measuring result is inherently secure against curious APs. Take RSS as an example, RSS values are measured at the user device by evaluating the signal strength from each AP without sending any information back. This kind of measurement leaks no information to the APs. If the communication during measurement is bilateral, colluding APs can obtain the full measuring result of the user and compute the location with the help of the localization server. Hence, it is crucial to make the measurements unlinkable or anonymous for measuring techniques involving bilateral communications.
4.1.2. Users
Users, the primary recipients of LBSs, are the central focus. The majority of LBSs are accessible via mobile devices. However, with the advancement of electronic technologies, wearable devices are also used for location applications and serving users. The data transmitted or received by users mostly contains sensitive information about their locations, jobs, habits, and POIs, emphasizing the need for robust privacy safeguards.
The privacy of the users is threatened in three different ways during the localization process. Firstly, the user may unintentionally reveal her location when she requests localization services, exposing her RSS/CSI information to attackers. Secondly, an untrusted party can capture the RSS/CSI measurements sent by the user, and use it to retrieve the user’s location from the server. Also, considering the server as an untrusted party, the server can determine the user’s location based on her RSS/CSI measurements.
The user itself can also act as a malicious party. Malicious users’ actions in indoor localization are twofold: 1) Malicious users can access unauthorized information on the server through security misconfiguration and hacking or capture other users’ data to track other users or sell the information for profit. The localization server and localization process should protect this data from being exposed; 2) Collision of malicious users is another threat. Peers subscribing to the same LBS can either collude to launch attacks, or one adversary can create fake peers to obtain the information they seek. For example by creating three fake anchor locations and using their corresponding distances to the target in an iterative trilateration based on a localization algorithm to obtain an inferred location (Liu et al., 2018).
4.1.3. Trusted Third Party (TTP)
TTP is an impartial and trustworthy entity that facilitates secure interactions and transactions between parties. In the context of privacy, a TTP is often involved in managing sensitive information, ensuring secure communication channels, and validating the identities of the parties involved. In indoor localization, it acts as a server dedicated to location privacy responsibilities, encompassing procedures such as anonymization and encryption, and is often managed by the LBS server or an arbitrarily trusted third party. Investigating the ownership and operational aspects of this server is crucial for understanding the trustworthiness of privacy-preserving measures.
As a trusted entity, it can access the transmitted data of users, which lacks adequate privacy protection. Consequently, an untrusted party may intercept and access the transmitted data by exploiting vulnerabilities within the TTP. TTPs themselves may also misuse the information they handle for unauthorized purposes, leading to privacy violations. This could involve the unauthorized sharing or selling of sensitive data.
4.1.4. LSP
The Location Service Provider (LSP) or LBS server is integral to responding to user inquiries and processing information for service provision. It contains important information about the building including the building map and location fingerprint database. It also receives and processes user information to provide location-based services to users. Therefore, protecting all these data is essential.
The inherent privacy risk of the LSB server lies where an untrusted party can gather real RSS/CSI measurements from various locations or generate an extensive set of artificial RSS/CSI measurements resembling those obtained through fuzzing. Subsequently, this entity can seek localization assistance from the server using the acquired measurements. By meticulously documenting pairs of RSS/CSI and their corresponding locations, the untrusted party can construct a fingerprint database resembling that maintained by the server. This deceptive replication of the server’s fingerprint database enables the untrusted entity to potentially manipulate location-based services, exploiting the similarity between the fabricated and authentic data to gain unauthorized access or compromise the integrity of the localization system.
From another point of view, the most formidable adversary within this context is the malicious localization server. In the typical functioning of a system, the ultimate determination of a device’s location heavily depends on the localization server. This pivotal role grants the server not only access to the location information of all mobile devices within its scope but also the capability to construct an approximate trajectory for a specific device through the analysis of multiple localization queries over time. The compromise of this server, whether through a sophisticated hack or other illicit means, opens the gateway to extracting significant information. The consequences extend beyond only location data, encompassing potentially sensitive details, user profiles, and any other information processed or stored within the server’s databases.
4.1.5. Content Provider
Some contents, including maps and POIs, are essential for LBS functionality. While some LBS providers employ a third party to deliver this service, others use their own content. For indoor localization, the content may include, the building map and semantic information of each section or room. Privacy of this content is also important as it has information about different parts of the buildings and the service provider would not like it to be publicly available, mostly for sensitive data. Consider a hospital where the service provider is reluctant for users to see the sensitive parts of the building, such as the electrical facilities, clean rooms, etc.
Content Providers, however, do not pose a privacy risk as their communication with the LSP is unilateral. However, there is a chance that an untrusted party adds unusable or inappropriate content to the system, which leads to poor servicing, and the users will stop using the system.
4.1.6. Data Publisher
After gathering user data, the collected dataset can be leveraged for various purposes. This data may include the building’s heat map for crowded locations or the frequently used trajectories. Investigating the utilization and potential re-identification risks of this published data is crucial for safeguarding user privacy.
The attacker can link the data of the published dataset with other public datasets to identify or re-identify users or other unauthorized activities. Similar to the Content Provider itself cannot introduce a privacy threat to the IPS, as the Data Publisher’s communication with the server is unilateral.
4.2. Data Structure-Driven Privacy Leakage
In this section, inherent privacy vulnerabilities in the data structure of the indoor location fingerprinting systems are introduced. The location data in LBSs varies in indoor environments due to the different methods used for indoor localization. Different from outdoor positioning systems, where the main data is the coordinates (e.g. longitude and latitude), the data for indoor environments employing the fingerprinting technique is various signal characteristics vectors. As shown in Fig. 5, the data in indoor location fingerprinting transmitted through the network as the tuple of $<ID,Time,location>$ (Wernke et al., 2014). The $ID$ and $Time$ data can be defined the same for indoor and outdoor localization, but the $location$ data has a different type. The terms are explained below:
Figure 5. Data structure in indoor location fingerprinting systems.
4.2.1. ID
Identity refers to a user’s name, email address, or any other characteristic that distinguishes one individual from another. Identifications in LBSs can either be consistent or inconsistent (Zheng and Zhou, 2011). Certain LBSs require constant user IDs, while others do not necessitate a user’s identity or a consistent one. For instance, users can employ LBSs under a fake name or in secret. The most common constant identity that LBSs require is an email address or a phone number. From another perspective, the ID can be hardcoded such as MAC-Address (or possibly IMEI). It can also be a costume ID given by software; for example, the session ID for data transmission or a self-defined ID is characterized by the application developer.
Identity-driven privacy leakage, commonly called deanonymization, aims to identify a user by gathering information. These leaks include:
  • Single identity leaks or personal identification assaults involve pinpointing a user’s identity through their indoor location, such as within a faculty building’s lab, or indirectly by narrowing down identification possibilities—such as inferring gender, educational level, or access to restricted areas. This approach allows for further investigation to potentially uncover someone’s identity (Krumm, 2007). For instance, a solitary presence in a room, like a teacher’s office, offers a higher chance of exact identification, while a crowded space, like a student-filled lab, lowers the probability of specific identification but doesn’t eliminate it.
  • Multiple identity information leaks, meeting disclosure leaks, or aggregated presence assaults try to determine the relationship between two or more people, or between people and an aggregated property, such as whether people meet at a particular time or the approximate number of persons visit a lab (Liu et al., 2018). As a case in point, when two or more people go to the same lab at specific time intervals, they are highly probably co-workers or friends, or they have a meeting at that time.
4.2.2. Location
The primary tool for determining the location of an object is spatial information. Coordinates, such as longitude and latitude, are employed for outdoor locations, while signal characteristics are used for indoor localization. Other information, like a store’s name, may also be associated with a location. In indoor fingerprinting localization, hence, the location data involve the characteristics of the existing signal in the indoor area, such as RSS or CSI when RF technologies are utilized for localization, as mentioned in Section 2. This location data is received by the user’s device, sent to the server for the calculation of the user’s location, and the estimated location is then sent back to the user. Location-driven privacy leakage refers to locating crucial areas, such as a person’s exact workplace in an indoor environment (Gu et al., 2016).
4.2.3. Time
Some LBSs integrate identification and location information, along with a timestamp for each location. Regarding time, data transmission can be categorized as either real-time or non-real-time. Certain LBSs forego real-time services, opting to collect data for later publishes. Nevertheless, the primary objective for most LBSs is to provide real-time services, each posing its own set of privacy issues. Real-time privacy protection is more challenging to achieve than non-real-time protection due to the need for significantly more scalable solutions. Moreover, given the unpredictable and dynamic motions of users, achieving global optimization with real-time information is exceedingly challenging (Zheng and Zhou, 2011). Two examples of LBSs that utilize real-time location data are navigation and augmented reality games.
Based on the combination of time and location data, the various types of spatial information can be broadly categorized into two groups: Single and Trajectory. Single location information refers to static spatial data that provides details about a specific point in the area. This category applies to situations where the focus is on a singular, fixed position, such as the layout of a room or the coordinates of a stationary object. In contrast, trajectory information involves the dynamic aspect of location data, capturing the movement and path of an object over time, such as tracking the trajectory of a person or an asset in indoor areas for navigation, asset management, or security purposes.
Time-driven privacy leakage itself may only lead to limited information, but when it is combined with location data it reveals sensitive information. This combination can be categorized as follows:
  • Location along with discrete time determines whether a user is present at a location at a specific time, where the presence or absence of the user may disclose information.
  • Location along with continuous time, or simply tracking, is to establish a user trace by putting all or part of the sequence of events together. Stalking is also the term used to describe this type of leak.
In summary, identity, location, and time trilogy constitute the sensitive data in IPS; thus, they are the main targets of the adversaries. This data is also considered dynamic, as it can change rapidly over time. Regardless of the data format and the dynamic nature of it, utilizing an LBS generates a substantial amount of location-related data. This data is highly correlated, as datasets of real-world places often exhibit significant coupling relations, and frequent connections between positions may provide more information than initially anticipated. However, an attempt to gain access to each identity, location, and time separately might not be helpful for the attackers. For example, accessing the identity by itself may not reveal anything meaningful, but when combined with the location and/or time, it can result in the disclosure of valuable information.
4.3. Inference-Driven Privacy Leakage
In indoor location fingerprinting, there are also some privacy leakages arising from the information inferred about the users (Hemkumar, 2024). Two critical aspects can be considered as the main dimensions that can be inferred and violate privacy in indoor location fingerprinting systems: identity inference and profiling completeness. Identity inference raises concerns about unauthorized tracking and monitoring, unveiling sensitive details through IPS. Simultaneously, profiling completeness, while enhancing user profiles, poses risks by holding sensitive information. These two privacy dimensions are explained below in the following subsections (Liu et al., 2018).
4.3.1. Identity inference
Identity inference from indoor location data introduces multifaceted privacy concerns. Location tracking through IPS not only unveils sensitive details about a user’s activities, habits, and interests but also poses the risk of unauthorized tracking and monitoring without user consent (Krumm, 2007). The identification capabilities of these systems, based on unique location patterns, raise ethical concerns related to targeted marketing, personal profiling, and potential discriminatory practices. Furthermore, the prospect of re-identification, where seemingly anonymous users can be linked to their identity by combining location data with external information, underscores the need for stringent privacy safeguards. Additionally, location data can lead to potential inference that extends beyond exact location disclosure, as frequent visits to specific places can reveal significant information, such as health conditions. Collectively, these aspects highlight the intricate challenges associated with preserving identity privacy in the context of location data and emphasize the importance of ethical considerations and robust protective measures.
4.3.2. Profiling completeness
Profiling completeness involves the frequent identification of Points of Interest (POIs) within a building based on user locations (Pombinho et al., 2011). Utilizing this information becomes crucial for gaining a deeper understanding of user behavior in real-world scenarios and predicting future actions. This data, commonly employed for marketing and surveillance purposes, possesses the capability to significantly enhance the comprehensiveness of profiles compared to basic data alone. The functionality of location data to seamlessly ”connect the dots” allows for the automated generation of profiles for individuals or organizations. IPS, pivotal for tracking a user’s position within a building, hold sensitive information. Unauthorized access to this information could lead to tracking and monitoring without user awareness or consent, echoing the challenges discussed in identity inference.
5. Attacks on Indoor Location Fingerprinting
An adversary attacks IPS aiming to collect the location information of users and use it for their benefit. This section discusses the attacks on indoor location fingerprinting systems, investigating the different aspects including adversary and attack models to obtain meaningful data about the user’s identity, location, and time.
5.1. Adversary Models
In general, user privacy means that no passive adversary (including a curious or malicious server) can determine the location of an honest user after intercepting all protocol messages. Server privacy, however, protects the server’s data from being compromised by malicious users using location queries (Yang and Järvinen, 2018). To discuss threat models on privacy preservation in indoor localization, we need to explain privacy assumptions on how entities can behave in IPS. Four assumptions are taken into consideration which are explained in the following subsection.
5.1.1. Fully trusted
In this setting, all parties involved are trusted to follow the protocol and not deviate from the agreed-upon behavior. This setting assumes that all participants will faithfully execute the protocol and will not attempt to gain unauthorized information or compromise the data being processed. This setting provides the lowest level of assurance in terms of privacy and requires a high degree of trust among the involved parties, which is very rare.
5.1.2. Semi-trusted (honest-but-curious)
In this setting, it is assumed that the parties involved will adhere to the protocol, but they may attempt to gain additional information from the protocol execution. In other words, while they will not intentionally deviate from the protocol, they may analyze the exchanged signals to infer additional information about the inputs or computations of other parties. Protocols designed for the Semi-trusted setting aim to prevent participants from learning more than they are supposed to, even if they try to gain information through analysis.
5.1.3. Fully malicious
In this setting, it is assumed that one or more parties may actively deviate from the protocol in an attempt to subvert its goals or compromise the privacy of the data. This setting represents the most adversarial scenario, where participants may collude, attempt to cheat, or launch attacks to learn more than they are allowed to or to undermine the privacy guarantees of the protocol. Protocols designed for a fully malicious setting typically incorporate stronger security mechanisms to withstand active attacks and attempts at manipulation.
5.1.4. Uunilateral-malicious
Introduced in (Yang and Järvinen, 2018), the unilateral-malicious setting is weaker than the fully malicious setting but stronger than the conventional Semi-trusted setting. We formulate the malicious behaviors specific to user sessions in the unilateral-malicious setting, such as manipulating Wi-Fi fingerprints and disclosing locations. We expect the server to act fairly honestly. In other words, while the server may be interested in learning the location of a user, it should run the protocol instance honestly to offer a high-quality service. Since a server’s malicious actions, such as dishonest executions, would be easily detected in practice and punished due to providing inaccurate location services, we can lower the security requirements for the server. In other words, if the service is poor, then users would likely just stop using the service and, consequently, make such an attack impossible. However, the server cannot easily identify a user’s malicious behaviors.
Understanding the difference between various settings lies in the level of trust and cooperation expected from the participants, as well as the defences and security measures needed to ensure privacy and integrity in the face of potential adversarial behaviors and threats.
Figure 6. Adversary and attack models overview in in indoor location fingerprinting systems.
5.2. Attack Models
In the context of computer security and privacy, an attack model refers to a representation of potential threats, adversaries, and their capabilities that a system or application may face. It helps security analysts, developers, and researchers understand and anticipate the possible ways in which a system might be compromised or exploited. Following the privacy assumptions, the attackers intend to seek data from IPS. As shown in Fig. 6, this section aims to explain the attack models for IPS, and how attackers exploit entities’ vulnerabilities and location fingerprinting data structure to obtain information about entities in IPS.
5.2.1. Location Privacy and Data Privacy Attacks
Location privacy revolves around the potential threats to an individual’s privacy resulting from the collection, use, and disclosure of their location information. Targeted advertising, monitoring daily habits, surveillance, stalking and harassment, discriminatory uses, and unauthorized access are examples of location privacy concerns. Indoor locations due to more precise positioning are more susceptible to privacy issues. Overall, location privacy is a subcategory of data privacy that is disparate in importance. By this means, from the perspective of the user, different locations have different privacy requirements. For instance, the majority of people worry significantly more about keeping their houses and workplaces private than they do about revealing where they visited in a shopping center. Despite differing levels of data sensitivity, the dangers associated with location privacy remain a substantial concern (Liu et al., 2018).
The major attack model on IPS is introduced in (Li et al., 2014), which is followed by most papers. As shown in Fig. 7, this attack model categorizes the threats into the Location Privacy Attacks and Data Privacy Attacks. In both of them, location and data can be attacked in two ways as follows:
  • Location Privacy Attack I: The attacker directly obtains the user’s location information from the query from the server.
  • Location Privacy Attack II: The attacker indirectly infers the user’s location information by accessing the user’s WiFi RSS vector.
  • Data Privacy Attack I: The attacker accesses the WiFi fingerprint database stored in the server.
  • Data Privacy Attack II: The attacker builds a WiFi fingerprint database, which is similar to the one stored in the LSP and the localization accuracy of using the first database is comparable with that of using the latter database.
The location privacy attack can be applied by passively capturing the signal on the communication channel between the user and the localization server, which can lead to important information. By either capturing user-to-server or server-to-user exchanged data, the attacker can get the location of the user. The data privacy attack can also occur by gaining direct access to data on the localization server via a hack or misconfiguration, or by querying patterned locations (e.g., querying for location on a patterned grid) to form the fingerprint database and consequently misuse it.
Figure 7. Location privacy and data privacy attacks in in indoor location fingerprinting systems.
Note that another categorization with the same principle is applied by using terms user-side privacy attack and server-side privacy attack (Yang and Järvinen, 2018). The localization server and the APs are regarded as adversaries for preserving the user’s location privacy. The user, on the other hand, is viewed as an adversary to protect the server’s database. The user must not reveal her location to untrusted parties, and the service provider should also keep its WiFi fingerprint database from unauthorized leaking. For example, a malicious user may download the database and sell it for profit.
5.2.2. Protocol-based attack
In this attack, the malicious actor uses cryptographic methods to extract data from encrypted data or discovers misconfiguration in data transmission protocols between user and server. Specifically in the chosen fingerprint attack (Yang and Järvinen, 2018) or known location attack (Yang and Järvinen, 2018), the user chooses special fingerprints, such as all-zeros or single-one fingerprints, to compromise the whole server’s database.
5.2.3. Inference attack
An inference attack is a type of attack in which an adversary deduces sensitive information about individuals or the system by analyzing seemingly innocuous or non-sensitive data. Inference attacks often exploit the relationships and patterns present in the available data to draw conclusions that may not be explicitly revealed. In the case of location-based services, inference attacks may attempt to deduce a person’s activities, habits, or even identity by analyzing patterns in the location data collected over time.
5.2.4. Semantic attack
In a semantic attack, the objective of the attacker is to obtain specific semantic details about an individual. Location data inherently contains semantic information, and the disclosure of such information can unveil personal details and behavioral tendencies of users. Within indoor location datasets, an adversary can deduce the purpose of a user’s visits to particular locations by monitoring their movements or extended stays (Fathalizadeh et al., 2022). For example, in a hospital, the adversary may discern whether the user is visiting a specific doctor, and in a shopping center, they can identify specific shops where the user lingers. Additionally, this information might reveal whether the user is searching for or purchasing particular products in a shopping center, consulting a specific specialist in a hospital, or meeting a particular professor at a university.
5.2.5. Probability-based attack
This attack type focuses on how the attacker would change his probabilistic belief about the sensitive information of a user after acquiring the captured data, rather than on what records, attributes, and tables the attacker may link to a target user. In general, this subset of privacy attacks seeks to achieve the uninformative principle, which aims to keep the gap between pre- and posterior beliefs as minimal as possible.
5.2.6. Context linking attack
Most location-based attacks make use of some context. Correlating metadata from different sources, such as social media profiles, public databases, or online activities, allows attackers to link contexts and uncover additional details about individuals. When launching a localization attack, it is simple to combine contextual knowledge with the observed location data to determine a target’s exact location. For instance, a personal context linking attack (Gruteser and Grunwald, 2003) may be used to remove all the unnecessary areas from a user’s location after reducing an obfuscated area to a specified spot. Contextual information can be used in conjunction with precise location data to carry out identity attacks. An attacker can assume that their target is or was in a special room of a hospital at a particular moment, for instance, if they know someone’s affiliation and discover it on a hospital room list.
5.2.7. Data linking attack
It is introduced in (Fathalizadeh et al., 2022) in whcih attacks on published indoor location datasets are considered twofold: identification/re-identification and Record/Attribute/Table linkage attacks. The attacker’s goal in an identification/re-identification attack is to identify/re-identify the victim in the whole dataset among all the other users that might be presented there. In an indoor location data set, an attacker with some background knowledge about an individual can identify whether he/she has been in a specific location. This is particularly important when spatial-temporal locations and trajectories are monitored by an adversary. A linkage attack occurs when an adversary can link a record owner (i) to a record in a published dataset, (ii) to a sensitive attribute in a revealed dataset, or (iii) to the published table of data by itself. These attacks are called record linkage, attribute linkage, and table linkage attacks, respectively. Linkage attacks are characterized by the adversary’s prior knowledge. In the case of record linkage and attribute linkage, the adversary knows that a specific individual’s data is presented in a dataset and he wants to learn about sensitive information. Table linkage attack, on the other hand, focuses on understanding whether a known individual’s information is available in the released dataset or not. The adversary could deduce sensitive information from the disclosed dataset based on the distribution of sensitive values in the group to which the individual belongs in an attribute linkage attack
6. Indoor Location Fingerprinting Privacy Preservation Mechanisms (ILFPPM)
This section reviews existing studies on ILFPPM. Four major privacy preservation techniques have been utilized for this purpose, including cryptographic methods, anonymization techniques, Differential Privacy (DP), and Federated Learning (FL). Each of these methods is discussed in the following sub-sections, accompanied by a brief explanation of the privacy method, necessary definitions, applicability of privacy methods to indoor settings, as well as their respective advantages and disadvantages.
Note that there are also passive solutions for localization (Schauer et al., 2016; Gschwandtner and Schindhelm, 2011), in which the fingerprinting dataset is available for the users and the localization process is done locally on the user’s device. In this method, however, the privacy of the user is preserved, but the server-side privacy, or in other words, server data privacy is not preserved, so it is not considered in this categorization.
6.1. Cryptographic Methods
Here, we briefly explain the cryptographic methods and then discuss the studies in which these methods are used in the concept of ILFPPM.
6.1.1. Cryptographic techniques
Definition 6.1 (Homomorphic encryption).
Homomorphic encryption is a cryptographic technique that enables computations to be executed on encrypted data directly, without the need for prior decryption. The outcomes of these computations remain encrypted, and upon decryption, yield results identical to those obtained had the operations been conducted on the original, unencrypted data. Two well-known homomorphic-based cryptosystems used in protecting the privacy of indoor location data are DGK (Damgard et al., 2008) and Paillier (Paillier, 1999). The DGK protocol was created by Damgård, Geisler, and Krøigaard in 2007 as an efficient solution to solving the millionaire’s problem. This protocol is a homomorphic cryptosystem dedicated to small plaintexts and fits nicely within the secure comparison protocol. The Paillier cryptosystem is a probabilistic asymmetric algorithm for public key cryptography. The scheme is an additive homomorphic cryptosystem; this means that, given only the public key and the encryption of $m_{1}$ and $m_{2}$, one can compute the encryption of $m_{1}+m_{2}$. It also allows the multiplication of a ciphertext by a plaintext number.
Figure 8. Homomorphic encryption vs SMPC in indoor location fingerprinting systems.
Definition 6.2 (Multi-Party Computations (MPC)).
Secure Multi-Party Computation (SMPC), or simply Multi-Party Computation (MPC), is a subfield within cryptography that aims to develop techniques enabling multiple parties to collaboratively compute a function using their inputs while maintaining the privacy of those inputs.
MPC protocols, especially in scenarios with unconditional private channels, rely on secret sharing. In secret-sharing-based (or Aritmethic secret sharing) methods (Shamir, 1979), the participating parties do not have specific roles, as seen in the creator-evaluator paradigm of Yao’s protocol. Instead, the data associated with each wire is shared among the parties, and a protocol is then employed to evaluate each gate. In this context, the function is defined as a ‘circuit’ over a finite field, unlike the binary circuits used in Yao’s protocol. Such circuits are referred to as ‘arithmetic circuits’ in the literature and consist of addition and multiplication ‘gates’ where the values being operated on are defined over a finite field.
Table 1. Studies on cryptography-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Adversary Model Attack Model (Li et al., 2014) 2014 kNN Paillier Simulated Private Coms. Overhead Loc. Error Fully Malicious Loc. & Data Priv. Attack (Wang et al., 2015) 2015 Fuzzy Logic Paillier CRAWDAD Loc. Error Semi-trusted Loc. Priv. Attack (Zhang et al., 2016) 2016 SVM Paillier Private Coms. Overhead Fully malicious Loc. & Data Priv. Attack (Yang and Järvinen, 2018) 2018 - Paillier Garbled Circuit - - Semi-trusted Loc. & Data Priv. Attack Protocol-based Attack (Yang and Järvinen, 2018) 2018 - Paillier Garbled Circuit Simulated - Unilateral-malicious Loc. & Data Priv. Attack Protocol-based Attack (Richter et al., 2018) 2018 kNN Paillier Garbled Circuit Private JUIndoorLoc Tampere Uni Loc. Error Floor Detection - Loc. Priv. Attack (Quijano and Akkaya, 2019) 2019 kNN Paillier DGK Algorithm Colombia Uni Execution Time Battery Life - - (Wang et al., 2019) 2019 kNN ABY Framework Arithmetic Sharing Boolean Sharing Yao’s Sharing Simulated Private same as (Li et al., 2014) Coms. Overhead Run Time Fully malicious Protocol-based attack (Järvinen et al., 2019) 2019 kNN Trusted Third Party Simulated Execution Time Energy Consumption Fully trusted Loc. & Data Priv. Attack Protocol-based Attack (Eshun and Palmieri, 2019) 2019 kNN Paillier Spatial Bloom - Coms. Overhead Semi-trusted Loc. Priv. Attack (Zhang et al., 2020) 2020 kNN Paillier Private Coms. Overhead Loc. Error Time Cost Fully malicious Loc. & Data Priv. Attack (Wu et al., 2020) 2020 kNN Expectation- maximization Bayes Network UJIIndoorLoc Coms. Overhead Loc. Error Time Cost Semi-trusted Loc. Priv. Attack (Nieminen and Järvinen, 2021) 2021 kNN Paillier One Time Pad Garbled Circuits Simulated Uni of Helsinki Coms. Overhead Loc. Error Time Cost Semi-trusted Loc. & Data Priv. Attack (van der Beets et al., 2022) 2022 kNN Arithmetic Sharing Delta Sharing Yao’s Sharing Simulated Coms. Overhead Run Time Semi-trusted - (Hu et al., 2022) 2022 MLE Paillier Private Coms. Overhead Loc. Error - - (Wang et al., 2024) 2024 kNN Inner Product Encryption Simulated JUIndoorLoc Loc. Error Coms. Overhead Time Cost Semi-trusted Data Priv. Attack
6.1.2. Cryptography-based ILFPPM
Cryptography-based ILFPPM uses encryption to protect users’ indoor locations. Encryption is the most popular mechanism, probably because it is a standard solution for securely transmitting data. Fig. 8 illustrates how can Homomorphic encryption and SMPC are utilized in indoor location fingerprinting systems.
Table 1 reports all the studies that employ Cryptography for ILFPPM. As can be seen, authors in (Li et al., 2014) present the first indoor location privacy preservation work by encrypting measured RSS. It also introduces the attack models for data privacy and location privacy which have been followed by other studies as well. However, it was found insecure in (Yang and Järvinen, 2018). Most papers’ localization process is done through kNN algorithm as can be seen in the Table; however, Support Vector Machine (SVM) (Zhang et al., 2016) and fuzzy logic (Wang et al., 2015) are also applied for probabilistic location estimation. As can be seen different cryptography-based approaches including Paillier cryptosystem, Garbled Circuits (GC), and DGK algorithm are applied in ILFPPMs. Most papers employ RSS values but the application of cryptography for CSI data is also introduced in (Wang et al., 2015). The most common metrics applied to evaluate the proposed schemes are computation and communication overhead and localization error. Jarvinen et al. propose PILOT (Järvinen et al., 2019), which was the first efficient solution for ILFPPM. PILOT employs well-known Secure Two-Party Computation (STPC) techniques and outsources the most involving computations to two semi-trusted servers to achieve practical performance. Zhang et al. (Zhang et al., 2016) shows a PPIL based on SVM and Paillier encryption, but (Yang and Järvinen, 2018, 2018) shows that it suffers from similar weaknesses. They discuss the state-of-the-art active attacks (theoretically) and server data privacy attacks. As a solution, they represent GC theoretically. In another work, Nieminen and Jarvinen (Nieminen and Järvinen, 2021) proposed a PPIL scheme in the user-server setting using Paillier encryption and GCs.
Cryptography methods operate without relying on TTP or outsourcing but they involve relatively expensive operations. Their scheme requires computationally heavy pre-computations in the setup phase for each localization query. These computations, especially on user devices, demand resources, posing a significant concern for devices with limited battery life.
6.2. Anonymization
Here, we briefly explain the anonymization techniques and then discuss the studies on anonymization-based ILFPPM.
(a) location
(b) trajectory
Figure 9. Examples of $3$-anonymity for location and trajectory privacy protection
6.2.1. Anonymization techniques
Data anonymization is a fundamental technique in data privacy and security, aiming to protect individuals’ sensitive information while still enabling the analysis and utilization of valuable datasets. Anonymizing data involves the transformation of personally identifiable information (PII) or sensitive attributes within a dataset in such a way that the identities of individuals are concealed, and their privacy is preserved. Several key anonymization techniques, including $k$-anonymity, $\ell$-diversity, and $t$-closeness, have been developed to achieve different levels of privacy protection and utility preservation in data sharing and analysis. These techniques are typically applied to quasi-identifiers, attributes that, when combined, can potentially lead to the identification of individuals.
Definition 6.3 ($k$-anonymity (Samarati and Sweeney, 1998)).
Given a dataset $\mathcal{D}$, a quasi-identifier $\mathcal{QI}$, and a parameter $k$, $\mathcal{D}$ satisfies $k$-anonymity if and only if, for each combination of quasi-identifier values in $\mathcal{QI}$, there exist at least $k$ records in $\mathcal{D}$ with identical quasi-identifier values.
Definition 6.4 ($\ell$-diversity (Machanavajjhala et al., 2006)).
Given a dataset $\mathcal{D}$, a quasi-identifier $\mathcal{QI}$, and a parameter $\ell$, $\mathcal{D}$ satisfies $\ell$-diversity if, for each combination of quasi-identifier values in $\mathcal{QI}$, there are at least $\ell$ ”well-represented” sensitive attribute values in the corresponding group of records.
Definition 6.5 ($t$-closeness (Li et al., 2007)).
Given a dataset $\mathcal{D}$, a quasi-identifier $\mathcal{QI}$, and a parameter $t$, $\mathcal{D}$ satisfies $t$-closeness if, for each combination of quasi-identifier values in $\mathcal{QI}$, the distribution of sensitive attribute values in the corresponding group of records is ”close” to the overall distribution of sensitive attribute values in the entire dataset.
Table 2. Studies on anonymization-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Adversary Model Attack Model (Kim et al., 2012) 2012 - $k$-anonymity Hierarchical graph Simulated Area of ASR No. of cells - Loc. Priv. Attack Inference Attack (Zhu and Zhai, 2014) 2014 kNN, SVM $k$-anonymity cloaking obfuscation Simulated Success Rate Area of ASR - Loc. Priv. Attack (Konstantinidis et al., 2015) 2015 - $k$-anonymity Bloom Filter Simulated CSUCY KIOSUCY Crawdad Energy Consumption No. of Messages Run Time Fully malicious Loc. & Data Priv. Attack Linking Attack (Kim and Li, 2016) 2016 kNN $k$-anonymity $\ell$-diversity Hierarchical graph Simulated Success Rate Area of ASR Response Time Fully trusted Loc. Priv. Attack Protocol-based Attack Context Linking Attack (Alikhani et al., 2018) 2018 NN Hilbert Curve $k$-anonymity Simulated Coms. Overhead Loc. Error - - (Sazdar et al., 2020b) 2020 kNN $k$-anonymity Randomization Permutation Simulated CRI Coms. Overhead Loc. Error Fully malicious Loc. Priv. Attack (Zhao et al., 2020) 2020 kNN $k$-anonymity Simulated Success Rate CR Area Coms. Overhead Loc. Error - Loc. Priv. Attack Context Linking Attack (Sazdar et al., 2020a) 2021 kNN $k$-anonymity Hilbert Curve Bloom Filter Simulated CRI CSUCY KIOSUCY PosData Loc. Error Semi-trusted Loc. priv. Attack (Fathalizadeh et al., 2022) 2022 - $k$-anonymity $\ell$-diversity $t$-closeness ($\alpha,k$)-anonymity $\delta$-presence Simulated CRI UJIIndoorLoc Spatial Loss RSS Loss NCP Semi-trusted Loc. & Data Priv. Attack Semantic Attack Data Linking Attack Probability-based Attack
6.2.2. Anonymization-based ILFPPM
Anonymization techniques can be employed in IPSs to protect the privacy of individuals while still enabling the provision of LBS. As illustrated in Fig. 9, these techniques can be applied to both location and trajectory data. The application of anonymization techniques for indoor localization was first introduced by authors in (Kim et al., 2012) using graph hierarchy, where nearby zones were mixed based on different algorithms. Due to $k$-anonymity’s vulnerability to identification and re-identification attacks, the proposed methods often combine with various approaches, such as Hilbert Curve, Bloom Filter, $\ell$-diversity, etc.
Table 2 reports the methods proposed for ILFPPMs based on anonymization, which indicates that the application of $k$-anonymity was first introduced by Hierarchical Graphs, followed by cloaking and obfuscation methods. The most widely used metric for analyzing these methods is the Success Rate to assess how often $k$-anonymity is satisfied. $k$-anonymity is also used for protecting the trajectory of the users inside the buildings (Sazdar et al., 2020b). The use of Hilbert Curve and Bloom Filter, both separately and in combination, is integrated with anonymization methods to enhance location privacy protection. In the latest work (Fathalizadeh et al., 2022) five methods are introduced, $k$-anonymity, $\ell$-diversity, $t$-closeness, ($\alpha,k$)-anonymity, $\delta$-presence, to defend against five major attacks.
6.3. Differential Privacy
Here, we provide a brief explanation of the Differential Privacy (DP) mechanism and various DP-based solutions and subsequently discuss the DP-based ILFPPM.
6.3.1. DP-based techniques
Definition 6.6 ($\epsilon$-differential privacy (Dwork et al., 2006)).
[…]
(2) $\frac{\Pr[\mathcal{A}(\mathcal{D})\in S]}{\Pr[\mathcal{A}(\mathcal{D}^{\prime})\in S]}\leq e^{\epsilon},$
where, the probability is computed over the randomness used by the algorithm.
Figure 10. CDP vs LDP in IPS
Definition 6.7 (Geo-indistinguishability (Andrés et al., 2013)).
When releasing aggregated information about a group of individuals, DP is an effective tool. However, it may not be suitable for scenarios involving a single individual, such as location privacy. DP prevents any change in an individual’s location from affecting the published output, making it challenging to convey meaningful information to the service provider. To address these limitations, Geo-Indistinguishability was introduced. Geo-Indistinguishability adds random noise to a user’s true location to prevent an adversary with background knowledge from accurately inferring the user’s location. This approach utilizes 2D Euclidean distance, suitable for localization purposes, whereas previous methods primarily focused on Hamming distance. Formally, for a randomized mechanism $\mathcal{K}$ that reports an obfuscated location $z$ instead of the user’s real location $x$, $\epsilon$-Geo-Indistinguishability is satisfied if, for any location ${x}^{\prime}$:
(3) $\mathcal{K}(x)(z)\leq e^{\epsilon d(x,{x}^{\prime})}\cdot\mathcal{K}({x}^{\prime})(z),$
where, $d$ represents the Euclidean distance between $x$ and ${x}^{\prime}$. Geo-Indistinguishability ensures that the probability of reporting point $z$ instead of other points $x$ and ${x}^{\prime}$ differs by at most $e^{-\epsilon d(x,{x}^{\prime})}$. This is achieved using a noise function, causing the probability of reporting a point around $z$ to decrease exponentially as the distance from the actual location $x$ increases. In a 1D linear space, this is accomplished using the Laplace distribution with the probability distribution function (PDF) $f_{X}(x)=\frac{\epsilon}{2}e^{-\epsilon\left|x-\mu\right|}$, where $\mu$ and $\epsilon$ are parameters in the Laplace distribution.
Table 3. Studies on different DP-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Adversary Model Attack Model (Li et al., 2016) 2016 kNN CDP Paillier Private Coms. Overhead Loc. Error Semi-trusted Loc. Priv. Attack (Sadhu et al., 2017) 2017 - LDP Two-step classifier Multi Building Rutgers Uni Loc. Error - - (Zhu et al., 2017) 2017 - LDP Delaunay triangulation Range generalization Simulated Loc. Error - - (Kim et al., 2018) 2018 - LDP Randomized Response Simulated Sangmyung Uni Error Rate Fully malicious Loc. Priv. Attack Probability-based Attack (Wang et al., 2018) 2018 - CDP AP fuzzification Finger Clustering Finger Permutation PosData Loc. Error Distance Error - Loc. & Data Priv. Attack Probability-based Attack (Zhao et al., 2018) 2018 - LDP Segmenting Data $k$-anonymity Private Coms. Overhead Loc. Error Semi-trusted Loc. Priv. Attack Context linking Attack Probability-based Attack Protocol-based Attack (Kim and Jang, 2019) 2019 - LDP Randomized Encoding Simulated Beijing Taxi Trajectory Loc. Error - - (Navidan et al., 2022) 2022 - LDP Local Hashing Unary Encoding Histogram Encoding Random response RAPPOR JUIndoorLoc CRI Loc. Error - Loc. & Data Priv. Attack (Zhang et al., 2022) 2022 FSELM LDP Edge–cloud collaboration Private Coms. Overhead Loc. Error Run Time Fully malicious Loc. Priv. Attack Probability-based Attack Inference Attack (Min et al., 2022) 2022 - GeoInd Laplace Mechanism Simulated Coms. Overhead Loc. Privacy QoS Loss Fully malicious Loc. Priv. Attack Probability-based Attack Inference Attack (Fathalizadeh et al., 2023) 2023 - GeoInd Laplace Mechanism Distance calculation RSS Generation Simulated JUIndoorLoc CRI Loc. Error Probabilistic QoS Loss Fully malicious Loc. Priv. Attack Probability-based Attack Context Linking Attack (Min et al., 2023) 2023 - GeoInd Reinforcement learning Simulated Coms. Overhead Loc. Privacy QoS Loss Fully malicious Inference Attack Semantic Attack Probability-based Attack (Hemkumar, 2024) 2024 - LDP GeoInd Geolife Gowalla QoS Loss Fully malicious Inference Attack Loc. Priv. Attack Probability-based Attack
6.3.2. DP-based ILFPPM
The privacy norm of DP is also utilized to protect privacy in indoor localization. The proposed methods lie on one of the Centralized Differential Privacy (CDP), Local Differential Privacy (LDP), or Geo-indistinguishability. CDP is a framework for protecting the privacy of individual data points in a centralized dataset. In this approach, a trusted server applies DP mechanisms to the entire dataset to ensure that the analysis or queries performed on the data do not reveal sensitive information about any individual data point. In LDP, on the other hand, individual data contributors apply privacy mechanisms to their data before sending it to a central server or aggregator. Each data point is perturbed or obfuscated in such a way that, when aggregated with other data points, the perturbed data still provides statistically valid information for analysis, but does not compromise the privacy of the individuals. This difference between CDP and LDP is plotted in Fig. 10. In addition, geo-indistinguishability is a DP-based technique used to protect the privacy of individuals’ locations by adding noise to local data, thus leading to obfuscating the real location.
Table 3 provides a list of ILFPPM based on DP. To illustrate, authors in (Li et al., 2016) utilize DP to inject noise into participants’ fingerprint data to protect location privacy. In contrast, the study conducted by (Zhu et al., 2017) extends DP and proposes a data obfuscation mechanism to generalize and specialize users’ fingerprints. Another study, as presented in (Sadhu et al., 2017), designs a localization platform relying on Wi-Fi RSS values and introduces data perturbation and randomization techniques grounded in DP. The aforementioned works concentrate on privacy preservation within specific types of indoor localization. The work described in (Zhao et al., 2018) introduces a privacy-preserving paradigm-driven framework for indoor localization utilizing local differential privacy. This approach is motivated by the observation that many IPS adhere to a common two-stage localization paradigm. Moreover, the study presented in (Kim and Jang, 2019) combines local differential privacy and optimal data encoding to disturb users’ data. Similarly, authors in (Wang et al., 2018) propose a DP-based privacy-preserving indoor localization scheme that introduces noise to the data, such as RSSI, CSI, and so forth, which increases the likelihood of significant localization errors and lowers the standard of localization services. The work in (Zhang et al., 2022) applies fusion of signal gathered from multiple wireless technologies (e.g., WiFi and BLE) with differentially private fingerprint fusion semi-supervised extreme learning machine for indoor localization in the edge computing, called Adp-FSELM. Another study introduces a novel privacy-aware framework for aggregating indoor location data employing the LDP technique. In this approach, the user location data is locally transformed on the user’s device and subsequently sent to the aggregator (Navidan et al., 2022).
Employing other DP-based solutions, the authors in (Min et al., 2022) initially introduce 3D Geo-indistinguishability for the application of indoor localization. This is followed by the addition of reinforcement learning for the sake of semantic location privacy in (Min et al., 2023). These papers assume that the user has access to her location and sends her location for obtaining services. From another perspective, (Fathalizadeh et al., 2023) introduces two methods for applying geo-indistinguishability based on RSS vectors without having access to the location.
Figure 11. An overview of FL in indoor location fingerprinting systems
6.4. Federated Learning
Here, we briefly explain the FL mechanism and then discuss the FL-based ILFPPM.
6.4.1. Federated Learning mechanism
Google debuted FL, a privacy-preserving distributed ML approach, in 2016 (McMahan et al., 2017). This method trains an algorithm on several decentralized edge devices or local servers without requiring them to exchange raw data, in contrast to traditional centralized machine learning techniques that load all local datasets onto a single server. Initially, a machine learning model is trained by a centralized server and this model is then sent to local devices that hold their own private data. On these respective devices, the local models are trained using their private data, as shown in Fig. 11. The training process typically involves using local computational resources and updates the model based on the local data. After training, the locally updated models send their changes (e.g., parameters or gradients) back to the centralized server, which aggregates these updates. Finally, the server updates the global model using an aggregation method such as FedAvg and FedAmp (Hsu et al., 2019; Wu et al., 2021a). FedAvg is one of the most common aggregation methods for FL, where the central server combines the gradients sent by the users using the below equation and updates the global model.
(4) $w_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}w^{k}_{t},$
where $n$ is the total number of data, $n_{k}$ is the number of samples for the $k$’th user, $w_{t+1}$ is the gradients of global model in the round $t+1$, $w^{k}_{t}$ is the gradients of $k$’th user’s local model in the round $t$. FL is especially useful in scenarios where data privacy or data locality is a concern. It allows for collaborative model training without the need to centralize sensitive data, thus mitigating privacy and security risks. By leveraging FL, systems can benefit from the collective intelligence of decentralized data sources while protecting user privacy and regulations.
Table 4. Studies on FL-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Attack Model (Liu et al., 2019) 2019 DNN AutoEncoder Private Coms. Overhead Loc. Error Loc. Priv. Attack (Li et al., 2020) 2020 SAEC FedAvg UJIIndoorLoc Loc. Error Loc. Priv. Attack (Ciftler et al., 2020) 2020 MLP FedAvg UJIIndoorLoc Loc. Error - (Wu et al., 2021b) 2021 MLP FedAvg FedAmp FedAmp-fusion UJIIndoorLoc Loc. Error - (Wu et al., 2021c) 2021 MLP FedAvg DP UJIIndoorLoc Guangzhou Xinguang shopping mall Loc. Error Noise Effect Inference Attack Loc. Priv. Attack (Tasbaz et al., 2022) 2022 kNN, MLP Random Forest FedAvg UJIIndoorLoc CRI Accuracy Run Time Loc. Priv. Attack Probability-based Attack (Etiabi et al., 2022) 2022 DNN Knowledge distillation Federated distillation Simulated UJIIndoorLoc Coms. Overhead Loc. Error - (Cheng et al., 2022) 2022 MLP FedAvg UJIIndoorLoc Accuracy Loc. Error - (Wu et al., 2022b) 2022 MLP FedAvg Knowledge distillation Mixture of Experts UJIIndoorLoc Guangzhou Xinguang shopping mall Loc. Error Inference Attack Loc. Priv. Attack (Nagia et al., 2022) 2022 CNN, LSTM FedProx Private Loc. Error - (Park et al., 2022) 2022 NN Monte Carlo Dropout UJIIndoorLoc Loc. Error Time Cost - (Wu et al., 2022a) 2022 GNN Federated Graph Learning Guangzhou Xinguang shopping mall Loc. Error - (Gao et al., 2023) 2023 CNN, DNN AutoEncoder UJIIndoorLoc Accuracy Loc. Error Coms. Overhead Inference Attack Loc. Priv. Attack (Kumar et al., 2023) 2023 CNN, LSTM - Private Accuracy - (Gufran and Pasricha, 2023) 2023 MLP AutoEncoder FedHIL Private Coms. Overhead Loc. Error Inference Attack (Guo et al., 2022) 2023 CNN Cloud Feature Extractor Homomorphic Encryption Simulated Coms. Overhead Loc. Error Inference Attack Probability-based Attack (Varma et al., 2023) 2023 kNN Discrete Coordinate Encryption Federated KNN Private Loc. Error Energy Consumption - (Yan et al., 2024) 2024 CNN Inner Production private Accuracy Loc. Error Data Priv. Attack
6.4.2. FL-based ILFPPM
As FL enables training a centralized model across decentralized devices or servers holding local data samples, without exchanging the data samples, it has been applied to IPS. IPS often involves collecting location data from various devices and using FL allows these devices to collaboratively learn a positioning model without sharing the raw location data. Accordingly, instead of sending all raw data to the central server, only model updates or aggregated information need to be communicated, thus reducing the amount of data transmitted over the network. Since indoor locations vary very widely, this approach is very useful when updating the model. So under the FL setting, indoor positioning models can be trained using data from various users in different environments without having to centralize that data. This helps protect the privacy of users while still allowing for model improvements and leads to a more accurate and robust IPS. Despite the benefits of applying FL in IPS, it also poses technical challenges including ensuring model convergence despite the heterogeneity of users’ data and addressing bias introduced by variations in local data. Table 4 reports ILFPPM studies in the FL setting. Most studies focus on the RSS data, but CSI is also taken into consideration in some papers. The proposed aggregation methods mostly are FedAvg and FedAmp, followed by widely used MLP and CNN learning algorithms.
6.5. Overall Comparison
This section discusses the metrics and datasets used in existing ILFPPM. Each ILFPPM considers various aspects of location privacy issues: cryptography-based ILFPPM mostly try to secure the connection channel between the server and the user and mostly lead to heavy computation and communication overheads. However, anonymization-based solutions implement lightweight protocols either on the device or server to obfuscate the localization mechanism. DP-based approaches encompass various solutions, including CDP, LDP, and geo-indistinguishability each addressing different aspects of the privacy norm within DP. Depending on the context, DP can be applied to both the server and the user. FL-based solutions aim to decentralize the learning mechanism for users to preserve privacy. Table 5 provides insights into the application areas and complexity levels of ILFPPM. Each privacy-preserving method, such as cryptographic methods employed primarily during transmission, anonymization implemented on both servers and devices, DP utilized across devices and servers, and FL mainly applied on devices. Additionally, the complexity levels in terms of computational requirements and expertise needed for the deployment of these mechanisms are reported in this Table.
Table 5. Overall comparison on ILFPPM.
ILFPPM Device/Transmission/Server Complexity
Cryptography Transmission High
Anonymization Device/Server Low
Differential Privacy Device/Server Medium
Federated Learning Device Medium
6.5.1. Evaluation metrics
Here, we discuss various metrics used in the different studies to evaluate the ILFPPM. All these metrics and their associated research are mentioned in Tables 1, 2, 3, 4. We categorize all metrics into the three main groups: localization accuracy, Quality of Service (QoS), and privacy metrics.
Localization Accuracy Metrics: The widely used metric in ILFPPM determines the localization accuracy. Location error metrics are crucial for evaluating the accuracy and precision of localization systems in various applications. These metrics are referred to by the following names: Loc Error, Distance Error, RMSE, MSE, MAE, Accuracy, Error Rate, Success Rate, Spatial Loss, RSS Loss, CR area, Area of ASR, floor detection, Kendall’s Tau distance, control value, count, and area. Location Error, Distance Error, Root Mean Square Error (RMSE), Mean Squared Error (MSE), and Mean Absolute Error (MAE) offer a measure of location error. Accuracy, Error Rate, and Success Rate represent the percentage of correctly localized instances, offering a straightforward measure of system performance. Spatial Loss and RSS Loss evaluate the discrepancy between predicted and actual locations within a spatial context, considering the geometric relationships between points. Additionally, metrics such as Cloaked Region (CR) Area, Area of ASR (Area under the Success Rate curve), and Floor Detection accuracy are tailored for specific localization contexts.
Quality of Service (QoS) Metrics: The second group is employed to assess the Quality of Service (QoS) of the ILFPPM. It includes various metrics, such as: QoS Loss, Communication Overhead, Time cost, Run Time, Execution time, Response time, Energy consumption, Entropy, number of messages, Noise Effect, Bandwidth cost, and battery life. QoS loss, measures the degradation in service quality concerning predetermined thresholds, indicating the system’s ability to meet desired performance levels. Computation and communication overhead quantify the additional computational or communication resources required to execute tasks, particularly relevant in cryptography-based approaches, where encryption adds computation and communication overheads without altering the data. Time cost metrics encompass various aspects such as execution time, reflecting the duration of task completion, and run time, which measures the duration of system operation, collectively shaping the system’s responsiveness and real-time performance. The number of messages exchanged during communication sessions provides insights into network traffic and resource utilization, influencing system scalability and efficiency. Noise effect metrics evaluate the impact of signal distortion or interference on data transmission and processing accuracy, essential for robustness in noisy environments.
Privacy Metrics: The last group consists of metrics related to privacy. Privacy metrics are essential for evaluating the effectiveness of privacy-preserving mechanisms in data handling and transmission. These metrics are referred to as Location Privacy, Probabilistic, and Normalized Central Penalty (NCP) in ILFPPM papers. Location privacy metrics evaluate the preservation of location privacy in spatial datasets, considering factors such as spatial granularity and proximity disclosure, crucial for protecting individuals’ location information. Probabilistic metrics for DP-based techniques assess the likelihood of privacy breaches in probabilistic databases, quantifying the probability of inferring sensitive information from query responses while preserving DP guarantees, crucial for ensuring privacy in data analysis and querying scenarios. NCP serves as a quantitative measure of privacy violation, indicating the extent to which sensitive information is disclosed in a dataset, with higher values reflecting increased privacy risks.
6.5.2. Datasets
Given the diversity of proposed methods and the fact that many of them may not be highly applicable to real-world datasets, most papers, even those intended for application to real-world datasets, employ their proposed methods in a simulated testbed. The open-source datasets widely used for indoor localization, even in non-privacy-focused papers, are UJIndoorLoc (Torres-Sospedra et al., 2014) and JUIndoorLoc (Roy et al., 2019). The CSUCY (Konstantinidis et al., 2015), KIOSUCI (Konstantinidis et al., 2015), CRAWDAD (Blinn et al., 2005), PosData (Wang et al., 2018), Geolife, and Gowalla (Hemkumar, 2024) are also public datasets that are used in ILFPPM papers. In some papers, the dataset’s location is not explicitly mentioned; instead, they mention the physical properties of the real-world dataset prepared for the proposed method. Some other datasets are Cyberspace Research Institute (CRI) (Moghtadaiee et al., 2019), Sangmyung University (Kim et al., 2018), Rutgers University (Sadhu et al., 2017), Tampere University (Lohan et al., 2017), Colombia University (Quijano and Akkaya, 2019), University of Helsinki (Nieminen and Järvinen, 2021) and Guangzhou Xinguang shopping mall (Chen et al., 2018) datasets.
7. Discussion & Future Research Directions
Indoor location privacy is a multifaceted challenge, given the diverse range of technologies and sensors used in positioning systems. To address these complexities and provide robust privacy protection while enabling the full potential of location-based services, future research can be summarized as follows:
Advanced hybrid privacy-preserving techniques: As every privacy-preserving technique has its own advantages and limitations, future research endeavors can be towards the development of advanced hybrid ILFPPM. These hybrid techniques aim to overcome the limitations of individual approaches, providing versatile solutions applicable across various sensor modalities and contextual scenarios. By combining different privacy-preserving strategies, these hybrids offer a comprehensive and adaptable framework for robust protection of user data in diverse indoor environments. In response to the evolving attack landscape, advanced hybrid methods can enhance privacy, defending against more potential attacks from breaching user privacy.
Context-aware privacy and usability: Future research could explore context-aware privacy mechanisms for indoor location data, incorporating user-friendly interfaces. This ensures that users can make informed decisions and manage privacy based on the sensitivity of location data in various contexts. The approach recognizes that privacy preservation is context-dependent and proposes techniques that dynamically adapt to different situations, such as stricter measures in healthcare and more leniency in retail settings. The aim is to develop adaptive privacy solutions capable of addressing concerns across diverse indoor applications.
Scalability: With the proliferation of IoT devices and the exponential growth of data, scalability challenges will need addressing. Advancements in FL can address scalability challenges by improving model architectures, communication efficiency, optimization algorithms, robustness against various attacks, and aggregation methods in large-scale IPS.
Ethical considerations and privacy standards: Future research on indoor location privacy needs to address ethical considerations related to location data use. Simultaneously, efforts should focus on developing privacy standards specifically designed for IPS. This integrated approach aims to ensure responsible and trustworthy deployment of LBS. In the evolving landscape of privacy regulations, such as the GDPR in Europe, new techniques must prioritize compliance, aligning privacy practices with legal requirements to establish a robust legal and ethical framework for indoor location data usage. Researchers should also work on the ethical implications of indoor location data, considering fairness, bias mitigation, and principled service deployment.
Privacy benchmarks and collaboration: Establishing privacy benchmarks and metrics is crucial for fostering collaboration among academia, industry, policymakers, and advocacy groups in the realm of ILFPPM. This collaborative ecosystem creates a shared framework for evaluating and advancing ILFPPM, incorporating research insights, practical implementations, informed policy perspectives, and user-centric advocacy. Privacy benchmarks act as a crucial link, bridging theoretical expertise with real-world applications to stimulate innovation. Policymakers can leverage these benchmarks to align regulations, ensuring a responsive and effective regulatory framework. User-centric principles, ethical considerations, and a commitment to continuous improvement serve as foundational elements in this collaborative effort, emphasizing responsible innovation and user trust as central tenets in the evolution of indoor location technologies.
Multimodal privacy integration: In the complex indoor environment, diverse sensor modalities are frequently employed, each possessing distinct strengths and limitations. Advanced approaches offer the capacity to seamlessly integrate data from these various sensors, culminating in comprehension of the indoor area, all while protecting user privacy. For example, the fusion of Wi-Fi and Bluetooth data with inputs from inertial sensors has the potential to significantly elevate not only the precision of indoor positioning but also the efficacy of privacy preservation mechanisms. This integrated approach empowers IPS to harness the complementary strengths of multiple sensor types, thereby enhancing both accuracy and privacy protection, ultimately affording users a more robust and private experience within indoor spaces.
8. Conclusion
This paper contributes a comprehensive overview of ILFPPM, presenting a foundational understanding of privacy preservation in indoor location fingerprinting. By addressing various dimensions, including definitions, applications, vulnerabilities, and protection mechanisms, we aim to provide more insights into the complexities of this dynamic domain. We introduce all possible origins of privacy leakages in indoor location fingerprinting. The novel categorization of adversary and attack models, along with the compilation of datasets and metrics, serves as a valuable resource for next empirical investigations. Through a detailed exploration of challenges, we point out the future research directions and opportunities in the field of privacy-preserving mechanisms for indoor location fingerprinting. With this extensive survey, we aim to establish a robust groundwork for forthcoming studies in this area.